4) Solve the initial value problem by Laplace Transform (10 marks) y - 2y +y = te y(0) = 1 %3D y(0) = 1 %3D
Answer #1

9-2y + y = det , & 10) =1, alo) = 1 Take lablace transform. s? YUS) - S-1-254689 + 9 +Y(S) = is? +1-95) VES) - S+1 = (3-1)2